OpenStudy (anonymous):

The area of a triangle must be at most 32.5sq in. The base is 13in and the height is x inches. Find the possible value of x. Use an inequality to solve the problem.

OpenStudy (anonymous):

okay so area of triangle = 1/2 * base * height we are given that area is 32.5sq in at most means that it is its maximum limit so 32.5 ≤ 1/2 * x* 13 65 ≤ 13x 5 ≤ x

OpenStudy (anonymous):

\[\le\] what is this mean

OpenStudy (anonymous):

what does that symbol mean \[\le\]

OpenStudy (anonymous):

it means that the area is equal or less than 1/2 * base*height